The Various Measures of Cost: Thirsty Thelma’s Lemonade Stand

Quantity

Of

lemonade

s


Fixed

Cost


Variable

Cost


Total

Cost


Average

Fixed

Cost


Average

Variable

cost


Average

Total

Cost


Marginal

Cost


0 $3 $0.00

1 3 0.3

2 3 0.8

3 3 1.5

4 3 2.4

5 3 3.5

6 3 4.8

7 3 6.3

8 3 8.0

9 3 9.9

10 3 12.0


Draw the graphs of TC, AVC, ATC, AFC and MC of the Thirsty Thelma’s Lemonade Stand(20)


1
Expert's answer
2021-02-01T12:43:44-0500

ANSWER

Let's write the formulas for unknown Total cost (denoted by TC), Average fixed cost (denoted by AFC), average variable cost (denoted by AVC), average total cost (denoted by ATC), marginal cost (denoted by MC):

"TC=FC+VC"

"AFC=FC\/Q"

"AVC=VC\/Q"

"ATC=(FC+VC)\/Q+AVC+ATC"

"MC(q)=TC(q)-TC(q-1)=VC(q)-VC(q-1)" .

Lets consider ane example of TC, AFC, AVC, ATC and MC when Q is equal to 3:

"TC(3)=3+1.5=4.5"

"AFC(3)=3\/3=1.0"

"AVC(3)=1.5\/3=0.5"

"ATC(3)=1.0+0.5=1.5"

"MC(3)=TC(3)-TC(2)=VC(3)-VC(2)=1.5-0.8=0.7"

The table that includes the calculated values is depicted in the picture 1. The graphs of TC, AVC, ATC, AFC and MC are depicted in the picture 2. The Ms Excel file is also attached.


Picture 1




Picture 2